OpenStudy (anonymous):

Identify the correct function for the following sequence. 18 , 15 , 12 , 9 , ...

OpenStudy (anonymous):

f(x) = 2x + 16 f(x) = x + 17 f(x) = -2x + 20 f(x) = -3x + 18